OpenStudy (anonymous):

Which set expresses the following numbers in scientific notation?: seven hundred-millionths, seven hundred million, and seven ten-thousandths.

OpenStudy (anonymous):

i think its the last one

OpenStudy (anonymous):

I don't know how to say it to you... Well, this is how I understood. 700 millionths = 7.0 × 10-4 700 million = 7.0 × 108 Seven ten-thousandths = 7.0 × 10-4 @myawilson Is that helpful? did u ask for the same ?